OpenStudy (anonymous):

Find the zeros of the function. State the multiplicity of multiple zeros. y = x(x - 1)3 x = 9 multiplicity 2 x = 0, x = 1 multiplicity 3 x = 0 multiplicity 2 x = 1 multiplicity 3

OpenStudy (solomonzelman):

If: \(\large\color{slate}{ 0=a\times b\times c \times d }\) then: \(\large\color{slate}{ a }\) , \(\large\color{slate}{ b }\) , \(\large\color{slate}{ c }\) , or \(\large\color{slate}{ d }\) is zero.

OpenStudy (solomonzelman):

you have a similar case: \(\large\color{slate}{ 0=x(x-1)^3 }\) (writing \(\large\color{slate}{ 0 }\) instead of \(\large\color{slate}{ f(x) }\) since you want to find the zeros)

OpenStudy (jdoe0001):

hint: \(\large { y = x(x - 1)^3 \\ \quad \\ 0= x(x - 1)^3\to \begin{cases} 0=x\\ 0=(x-1)^3\to 0=(x-1)(x-1)(x-1) \end{cases} }\)

OpenStudy (solomonzelman):

you have a similar case: \(\large\color{slate}{ 0=x(x-1)^3 }\) (writing \(\large\color{slate}{ 0 }\) instead of \(\large\color{slate}{ f(x) }\) since you want to find the zeros)

OpenStudy (anonymous):

so D i think

OpenStudy (solomonzelman):

what are the zeros, of the functions? which of the options has them/it listed?

OpenStudy (anonymous):

i forgot about the other x gah but this helped me a lot thank you @SolomonZelman

OpenStudy (solomonzelman):

Oh, surely:) I can see you got this this.... This is what I am here for; not a problem!